Preliminary plans are under way for the construction of a new stadium for a major league baseball team. City officials have questioned the number of profitability of the luxury corporate boxes planned for the upper deck of the stadium. Corporations and deleted individuals may buy the boxes for $300,000 each. The fixed construction cost for the upper-deck area is estimated to be $4,500,000, with a variable cost of $150,000 for each box constructed. A) Write an expression for total cost B) Write an expression for total revenue C) Write an expression for total profit
Answer #1

A)

Total cost = total Fixed cost + total Variable cost

Total fixed cost = $4,500,000

Total Variable cost = unit variable cost*Quantity

Unit variable cost = $150,000

Quantity =Q

Total cost = 4,500,000+(150,000*Q)

B)

Total revenue = Price *Quantity

Price = 300,000

Total revenue = 300,000 *Q

C) Total profit = Total revenue - Total cost = 300,000 *Q - (4,500,000+(150,000*Q) )

Total profit = 150,000*Q - 4,500,000